# Optimistic Online Non-stochastic Control via FTRL

### v1

@@ IV Online control with Optimistic FTRL
$f_{t}^{(i)}(M)\doteq\begin{cases}\langle{\boldsymbol{\alpha}_{t}},{\boldsymbol{\psi}^{i-1}_{t}(M)}\rangle&\text{if }i\geq 1\\ \langle{\boldsymbol{\beta}_{t}},{\boldsymbol{\psi}_{t}(M)}\rangle&\text{if }i=0\end{cases}$ (7)
@@
$\boldsymbol{\psi}_{t}(M)$ $\doteq M\overline{\boldsymbol{w}}_{t-1},$ (9)
@@
$G_{t}^{(i)}=\begin{cases}B^{\top}(A^{i-1})^{\top}\boldsymbol{\alpha}_{t}\overline{\boldsymbol{w}}_{t-i-2}^{\top}&\text{if }i\geq 1\\ \boldsymbol{\beta}_{t}\overline{\boldsymbol{w}}_{t-1}^{\top}&\text{if }i=0\end{cases}$ (10)
@@
$\|G_{t}^{(i)}\|\leq\begin{cases}\alpha\kappa_{B}pw(1-\delta)^{i-1}&\doteq g^{(i)}\quad\text{if }i\geq 1\\ \beta pw&\doteq g^{(0)}\quad\text{if }i=0\end{cases}$ (11)
using that for any matricies $A$,$B$, and vector $\boldsymbol{w}$ $\|AB\|=\|A\|\|B\|$ and $\|A\boldsymbol{w}\|\leq\|A\|_{\text{op}}\|\boldsymbol{w}\|$. We also define the prediction $\tilde{G}_{t}^{(i)}$, which we can construct by plugging the oracle’s output in (10), and hence we have the partial prediction error:
@@
We highlight that from (11), the magnitude of prediction error decrease exponentially with $i$: $\Delta_{t}^{(i)}\!\!\propto\!(1-\delta)^{i}\leq e^{-i}$. We refer to $i$ therefore as the attenuation level.
Similarly, we define ${G}_{t}$ as the matrix of partials of $F_{t}(\cdot)$, with its prediction as ${\tilde{G}}_{t}$, Finally, we define the hybrid hint matrix $H_{t}$, which aims to approximate the sum $G_{t-d:t}$
$\!\!\!\!H_{t}\!\doteq\!\bigg{\{}\underbrace{\!\sum_{i=0}^{d-1}\sum_{j=0}^{d-i-1}\!\!G_{t-d+i+j}^{(j)}}_{\text{available at $t$}}\ \!\!+\!\!\!\!\underbrace{\sum_{j=d-i}^{d}\!\!\tilde{G}_{t-d+i+j}^{(j)}\!\bigg{\}}\mathds{1}_{d\geq 1}+\!\tilde{G}_{t}}_{\text{future predictions}}$ (13)
We denote by $\Delta_{t}$ the prediction error $\Delta_{t}\doteq\|G_{t-d:t}-H_{t}\|$. Due to the definition of $F_{t}(\cdot)$, certain elements of the summands (in $G_{t-d:t}$) are partially observed by $t$ and are hence directly used in constructing $H_{t}$. The remaining elements in the sum are obtained from the prediction oracle.
Figure 1: Our methodology in designing OptFTRL-C. Up: The NSC to OCO-M reduction, with parameter $d=2$ (Sec. IV-A). Down: an equivalent delayed OCO formulation, obtained via rearrangement, which we append with an oracle (Sec. IV-B).
@@
0:  Policy parameters $M_{t}$ at each slot $t=1,\ldots,T$.
@@
6:     Receive future predictions $\{c_{\tau}(\cdot,\cdot)\}_{\tau=t}^{t+d}$
@@
The steps of the OptFTRL-C routine are outlined in Algorithm 1. OptFTRL-C first executes an action $\boldsymbol{u}_{t}$ (line $2$). Then, the cost function is revealed, completing the necessary information to compute $G_{t-d}$ (line $3$, recall that all the costs from $t-d,\dots,t$ are required to know $G_{t-d}$). The system then transitions to state $\vec{x}_{t+1}$, effectively revealing the disturbance vector $\vec{w}_{t}$ (line $4$). At this point, we can calculate $\Delta_{t-d}$ (since we know the ground truth $G_{t-d}$) and update the strong convexity parameter accordingly (line $5$). Then, the oracle forecasts the next $d$ costs functions (line $6$), enabling us to construct the hybrid hint matrix (line $7$). Finally, the next action $\boldsymbol{u}_{t+1}$ is committed through updating the policy parameters (line $8$). The regret of this OptFTRL-C is characterized in the following theorem:
@@
Let $(A,B)$ be an LTI system with the memory parameter $d$ defined as in Lemma 1. Let $\{c_{t}(\cdot,\cdot)\}_{t=1}^{T}$, $\{\boldsymbol{w}\}_{t=1}^{T}$ be any sequence of costs and disturbances, respectively. Let $\Delta^{(i)}_{t}$ be the prediction error for these sequences at time $t$ with attenuation level $i$, as defined in (12). Then, under Assumptions $1$ and $2$, algorithm OptFTRL-C produces actions $\{\boldsymbol{u}_{t}\}_{t=1}^{T}$ such that for all $T$, the following holds:
$\mathcal{R}_{T}=\mathcal{O}\left(\!\sqrt{\sum_{t=1}^{T}\!\bigg{(}\!\sum_{i=0}^{d}\sum_{j=i}^{d}\!\Delta^{\!(d-j+i)}_{t+i}\!\!\bigg{)}^{\!\!2}}\ \right).$
Discussion. OptFTRL-C achieves the sought-after accuracy-modulated regret bound that holds for systems with memory. It strictly generalizes previous optimistic online learning bounds by incorporating memory (hence states), and strictly generalizes previous online non-stochastic control bounds by handling predictions of unknown quality. Namely, OptFTRL-C’s bound has the following characteristics.
@@
$\leq 2\beta pw+\frac{2(1+\delta)\alpha\kappa_{B}pw}{\delta^{2}}\doteq m$
where the first equality follows from the triangular inequality and (11), and in the last inequality we used $\sum_{i=0}^{\infty}\!i(1\!-\!\delta)^{i}\!\leq\!\nicefrac{{1\!-\!\delta}}{{\delta^{2}}}$ and $\sum_{i=0}^{\infty}(1\!-\!\delta)^{i}\!\!\leq\!\nicefrac{{1}}{{\delta}}$. Hence, the regret becomes $\mathcal{O}(m\sqrt{T})$, which is order-optimal in $T$ [36, Sec 5.1], achieving the optimistic premise.
Memory-commensurate: Apart from prediction adaptivity, OptFTRL-C’s performance is interpretable with respect to the spectrum of stateless to stateful systems. Consider the stateless case ($\boldsymbol{x}_{t}=\boldsymbol{0},\forall t)$. Hence, $c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\langle{\boldsymbol{\beta}_{t}},{\boldsymbol{u_{t}}}\rangle$. In this case, OptFTRL-C requires predictions only for the next step (as per (13)), The resulting bound becomes $\mathcal{O}((\sum_{t=1}^{T}\Delta^{0}_{t})^{\nicefrac{{1}}{{2}}})=\mathcal{O}((\sum_{t}\|\boldsymbol{\beta}_{t}-\tilde{\boldsymbol{\beta}}\|_{t})^{\nicefrac{{1}}{{2}}})$, recovering the optimistic bound for stateless online learning [10].
@@
Now that we have described OptFTRL-C and characterized its policy regret, we present the main tools necessary to prove Theorem 1. Our proof is structured into two primary parts. First, we demonstrate that the regret in linearized NSC is a specific instance within the OCO-M framework, achieved through a particular selection of separable functions (sub-section IV-A, visualized in the upper timeline of Fig. 1). Second, we establish that the regret of OCO-M with separable functions is in turn a particular case within the Delayed OCO (OCO-D) framework with a specific structure of delay (sub-section IV-B, visualized in the lower timeline of Fig. 1). While the first part is fairly standard in NSC, we do not reduce its resulting OCO-M instance to standard OCO, but to OCO-D instead.
@@
Given $d\geq\frac{1}{\delta}\log(\frac{z}{\delta\epsilon}T)$, $z\doteq w(\kappa_{B}\kappa_{M}p+1)$. Then for any $\epsilon>0$, the following holds
$\sum_{t=1}^{T}\bigg{|}c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})-\sum_{i=0}^{d}f_{t}^{(i)}(M_{t-i})\bigg{|}\leq l\epsilon.$
@@
For now, assume that $\|\hat{\boldsymbol{x}_{t}}-\boldsymbol{x}_{t}\|\leq\frac{\epsilon}{T}$. Then, by Lipschitzness:
$\sum_{t=1}^{T}|c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})-c_{t}(\boldsymbol{\hat{x}}_{t},\boldsymbol{u}_{t})|<l\epsilon.$ (18)
However, $c_{t}(\boldsymbol{\hat{x}}_{t},\boldsymbol{u}_{t})$ can be written in terms of $f_{t}^{(i)}(\cdot)$ using Assumption $1$ and the definitions in (8) and (9) :
@@
It remains to show that $\|\boldsymbol{x}_{t}-\hat{\boldsymbol{x}_{t}}\|\leq\frac{\epsilon}{T}$. From (5) and (17), we have $\|\hat{\boldsymbol{x}_{t}}-\boldsymbol{x}_{t}\|\leq$
$\|\sum_{i=d}^{t-1}A^{i}\big{(}BM_{t-i-1}\overline{\boldsymbol{w}}_{t-i-2}+\boldsymbol{w}_{t-i-1}\!\big{)}\|$
$\leq\sum_{i=d}^{t-1}(\|A^{i}BM\|_{op}\|\|\boldsymbol{\overline{w}}_{t-i-2}\|+\|A^{i}\|_{op}\|\boldsymbol{w}_{t-i-1}\|)$
@@
where $(a)$ follow from the sub-multipilicitive property of $\|\cdot\|_{\text{op}}$, and $\|B\|_{\text{op}}\ \leq\|B\|$, and $(b)$ from $1-x\leq e^{-x}$. Substituting $d$, makes the last term $\frac{\epsilon}{T}$. ∎
The closeness of $\boldsymbol{x}_{t}$ and $\boldsymbol{\hat{x}}_{t}$ is a known fact in NSC, and it is due to the (strong) stability assumption. Lastly, we note that it is possible to set $d$ adaptively without knowledge of $T$ by using $d=\frac{1}{\delta}\log(\frac{zt}{\delta\epsilon})$. [^5] This would result in a sum of the form $\sum_{t}\nicefrac{{1}}{{t}}\leq\log(T)$. We leave this part for the extended version and, inline with previous works such as [15], consider here the $d$ is sufficiently large to have the approximation error $\epsilon$ as a negligible constant w.r.t. $T$.
@@
Let $f_{t}(M_{0},\dots,M_{d})$ be a separable function: $f_{t}(M_{0},\dots,M_{d})=\sum_{i=0}^{d}f_{t,i}(M_{i})$, $f_{t,i}(M):\mathbb{R}^{(d_{u}\times d_{x}p)}\mapsto\mathbb{R}$ Let $\mathcal{A}$ be an online learning algorithm whose decisions $M_{t+1}$ depend on the history set $\mathbb{H}_{t}\doteq\{\cup_{i=0}^{d}f_{\tau,i}(\cdot)\}_{\tau=1}^{t}$. Define $J_{t}(M)\doteq\sum_{i=0}^{d}f_{t+i,i}(M)$, and the history set w.r.t. $J_{t}(\cdot)$ as $\mathbb{H}^{J}_{t}$ Then,
@@
(20) means that the accumulated cost, with memory, is equivalent to that of the memoryless function $J_{t}(\cdot)$. However, from (19) $\mathcal{A}$ has delayed feedback w.r.t. $J_{t}(\cdot)$; when deciding $M_{t+1}$, feedback up to only $t-d$ is available.
@@
the first part is immediate from the definition of $J_{\tau}(\cdot)$; any $J_{\tau}(\cdot)$ with $\tau>t-d$ would require a function that is not in the original history set $\mathbb{H}_{t}$.
@@
Where the first equality holds by separability, the second by shifting the sum index and using the convention $f_{\tau}(M_{t<1})\doteq 0,\forall\tau$, the third by using $f_{t>T}(\cdot)\doteq 0$.∎
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{t=1}^{T}\big{(}\sum_{i=0}^{d}f_{t}^{(i)}(M_{t-i})-\sum_{i=0}^{d}f_{t}^{(i)}(M^{\star})\big{)}+2l\epsilon$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{t=1}^{T}f_{t}(M_{t-d},\dots,M_{t})-f_{t}(M^{\star},\dots,M^{\star})$ (22)
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{t=1}^{T}F_{t}(M_{t})-F_{t}(M^{\star})\doteq\mathcal{R}_{T}^{F},$ (23)
where $(a)$ follows by Lemma 1, which gives both an upper bound on the learner cost and lower bound on the benchmark’s cost, $(b)$ by writing the sum of $f_{t}^{(i)}(\cdot)$ as a single function with memory, and $(c)$ by Lemma 2, with $f_{t,i}(M_{t-i})=f_{t}^{(i)}(M_{t-i})$, and hence, $J_{t}(M)=F_{t}(M)$. To bound the delayed feedback regret $\mathcal{R}_{T}^{F}$, we use [16, Thm. 11], which we restate below using the notation of this paper: [^6] Mapping their notation to ours, we have $\boldsymbol{g}_{t-D:t}=G_{t-d:t}$, $\boldsymbol{h}_{t}=H_{t}$, $\boldsymbol{g}_{t-D:t}-h_{t}=\Delta_{t}$, $\boldsymbol{a}_{t,F}=2\kappa_{M}\Delta_{t}$, $\boldsymbol{b}_{t,F}=\nicefrac{{1}}{{2}}\Delta^{2}_{t}$, and $\alpha=\kappa_{M}^{2}$.
Theorem 2 ([16, Thm. 11])
Let $\lambda_{t}$ be non-decreasing on $t$ defined as in (16). Let $r(\cdot)$ be $\lambda_{t}$ strongly convex regularizer. Then, the sum in (23), can be bounded as:
$\mathcal{R}_{T}^{F}\leq 8\kappa_{M}\max_{t\in[T]}\Delta_{t-d:t-1}+2\sqrt{5}\kappa_{M}\sqrt{\sum_{t=1}^{T}\Delta_{t}^{2}}.$
@@
Now, note that the sum $G_{t-d:t}$ can be written as:
@@
hence we get that $\Delta_{t}=\|G_{t-d:t}-H_{t}\|\leq$
$\sum_{i=0}^{d-1}\sum_{j=d-i}^{d}\|{G}^{(j)}_{t-d+i+j}-{\tilde{G}}^{(j)}_{t-d+i+j}\|+\sum_{j=0}^{d}\Delta^{(j)}_{t+j}$ (25)
$=\sum_{i=0}^{d}\sum_{j=d-i}^{d}\Delta_{t-d+i+j}^{(j)}=\sum_{i=0}^{d}\sum_{j=0}^{i}\Delta_{t+j}^{(j+d-i)}$ (26)
$=\sum_{i=0}^{d}\sum_{j=i}^{d}\Delta_{t+i}^{(d-j+i)}$ (27)
substituting gives the result. ∎

### v2

@@ IV Online control with Optimistic FTRL
$f_{t}^{(i)}(M)\doteq\begin{cases}\langle{\boldsymbol{\alpha}_{t}},{\boldsymbol{\psi}^{i-1}_{t}(M)}\rangle&\text{if }i\geq 1,\\ \langle{\boldsymbol{\beta}_{t}},{\boldsymbol{\psi}_{t}(M)}\rangle&\text{if }i=0,\end{cases}$ (7)
@@
$\boldsymbol{\psi}_{t}(M)$ $\doteq M\overline{\boldsymbol{w}}_{t-1}.$ (9)
@@
$G_{t}^{(i)}=\begin{cases}B^{\top}(A^{i-1})^{\top}\boldsymbol{\alpha}_{t}\overline{\boldsymbol{w}}_{t-i-2}^{\top}&\text{if }i\geq 1,\\ \boldsymbol{\beta}_{t}\overline{\boldsymbol{w}}_{t-1}^{\top}&\text{if }i=0.\end{cases}$ (10)
@@
$\|G_{t}^{(i)}\|\leq\begin{cases}\alpha\kappa_{B}pw(1-\delta)^{i-1}&\doteq g^{(i)}\quad\text{if }i\geq 1,\\ \beta pw&\doteq g^{(0)}\quad\text{if }i=0,\end{cases}$ (11)
using that for any matrices $A$,$B$, and vector $\boldsymbol{w}$ $\|AB\|=\|A\|\|B\|$ and $\|A\boldsymbol{w}\|\leq\|A\|_{\text{op}}\|\boldsymbol{w}\|$. We also define the prediction $\tilde{G}_{t}^{(i)}$, which we can construct by plugging the oracle’s output in (10), and hence we have the partial prediction error:
@@
We highlight that the prediction error magnitude decreases exponentially with $i$: $\Delta_{t}^{(i)}\!\!\propto\!(1-\delta)^{i}\leq e^{-i}$. This follows from (11) and the fact that predicting $\tilde{G}_{t}^{(i)},i\geq 1$ amounts to predicting $\boldsymbol{\alpha}_{t}$ and attenuating it by a multiplication with $A^{i-1}$. We refer to $i$ therefore as the attenuation level.
Similarly, we define ${G}_{t}$ as the matrix of partial derivatives of $F_{t}(\cdot)$, and ${\tilde{G}}_{t}$ as its prediction. Lastly, we define the hybrid hint matrix $H_{t}$, which aims to approximate the sum $G_{t-d:t}$
$\!\!\!\!H_{t}\!\doteq\!\underbrace{\!\sum_{i=0}^{d-1}\!\bigg{(}\sum_{j=0}^{\begin{subarray}{c}d-i-1\end{subarray}}\!\!G_{t-d+i+j}^{(j)}}_{\text{observed at $t$}}\ \!\!+\!\!\!\!\underbrace{\sum_{j^{\prime}=d-i}^{d}\!\!\!\tilde{G}_{t-d+i+j^{\prime}}^{(j^{\prime})}\!\bigg{)}\mathds{1}_{d\geq 1}\!+\!\tilde{G}_{t}}_{\text{future predictions}},$ (13)
and we denote by $\Delta_{t}$ the prediction error $\Delta_{t}\doteq\|G_{t-d:t}-H_{t}\|$. Due to the definition of $F_{t}(\cdot)$, certain elements of the summands (in $G_{t-d:t}$) are partially observed at $t$ and are hence directly used in constructing $H_{t}$. The remaining elements in the sum are obtained from the prediction oracle.
Figure 1: The methodology of designing OptFTRL-C. Up: The NSC to OCO-M reduction, with parameter $d=2$ (Sec. IV-A). Down: an equivalent delayed OCO formulation, obtained via rearrangement, which we append with an oracle (Sec. IV-B).
@@
0:  Actions $\boldsymbol{u}_{t}$ at each slot $t=1,\ldots,T$.
@@
6:     Receive future predictions $\{c_{\tau}(\cdot,\cdot)\}_{\tau=t+1}^{t+d}$
@@
The steps of the OptFTRL-C routine are outlined in Algorithm 1. OptFTRL-C first executes an action $\boldsymbol{u}_{t}$ (line $2$). Then, the cost function is revealed, completing the necessary information to compute $G_{t-d}$ (line $3$, recall that all the costs from $t-d,\dots,t$ are required to know $G_{t-d}$). The system then transitions to state $\vec{x}_{t+1}$, effectively revealing the disturbance vector $\vec{w}_{t}$ (line $4$). At this point, we can calculate $\Delta_{t-d}$ (since we know the ground truth $G_{t-d}$) and update the strong convexity parameter accordingly (line $5$). Then, the oracle forecasts the next $d$ costs functions (line $6$), enabling us to construct the hybrid hint matrix (line $7$). Finally, the next action $\boldsymbol{u}_{t+1}$ is committed through updating the policy parameters (line $8$). The regret of this routine is characterized in the following theorem:
@@
Let $(A,B)$ be an LTI system, and $\{c_{t}(\cdot,\cdot)\}_{t=1}^{T}$, $\{\boldsymbol{w}\}_{t=1}^{T}$ be any sequence of stage costs and disturbances, respectively. Let $\Delta^{(i)}_{t}$ be the prediction error at $t$ with attenuation level $i$, as defined in (12). Then, with memory parameter $d$ defined as in Lemma 1, and under Assumptions $1$ and $2$, algorithm OptFTRL-C produces actions $\{\boldsymbol{u}_{t}\}_{t=1}^{T}$ such that for all $T$, the following holds:
$\mathcal{R}_{T}=\mathcal{O}\left(\!\sqrt{\sum_{t=1}^{T}\!\bigg{(}\!\sum_{i=0}^{d}\sum_{j=i}^{d}\!\Delta^{\!(d-j+i)}_{t+i}\!\!\bigg{)}^{\!\!2}}\ \right).\vspace{2mm}$
Discussion. OptFTRL-C achieves the sought-after accuracy-modulated bound that holds for systems with memory. It generalizes previous optimistic online learning bounds by incorporating memory (hence states), and generalizes previous online non-stochastic control bounds by handling predictions of unknown quality. Namely, OptFTRL-C’s bound has the following characteristics.
@@
$\leq 2\beta pw+\frac{2\alpha\kappa_{B}pw}{\delta^{2}}\doteq m$
where the first equality follows from the triangular inequality and (11), and in the last inequality we used $\sum_{i=0}^{\infty}\!i(1\!-\!\delta)^{i}\!\leq\!\nicefrac{{1\!-\!\delta}}{{\delta^{2}}}$ and $\sum_{i=0}^{\infty}(1\!-\!\delta)^{i}\!\!\leq\!\nicefrac{{1}}{{\delta}}$. Hence, the regret becomes $\mathcal{O}(m\sqrt{T})$, which is order-optimal in $T$ [35, Sec 5.1], achieving the optimistic premise.
Memory-commensurate: Apart from prediction adaptivity, OptFTRL-C’s performance is interpretable with respect to the spectrum of stateless to stateful systems. Consider the stateless case ($\boldsymbol{x}_{t}=\boldsymbol{0},\forall t)$. Hence, $c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\langle{\boldsymbol{\beta}_{t}},{\boldsymbol{u_{t}}}\rangle$. In this case, OptFTRL-C requires predictions only for the next step (as per (13)), The resulting bound becomes $\mathcal{O}((\sum_{t=1}^{T}\Delta^{0}_{t})^{\nicefrac{{1}}{{2}}})=\mathcal{O}((\sum_{t}\|\boldsymbol{\beta}_{t}-\tilde{\boldsymbol{\beta}}\|)^{\nicefrac{{1}}{{2}}})$, recovering the optimistic bound for stateless online learning [10].
@@
We now present the tools to prove Theorem 1. Our proof is structured into two primary parts. First, we demonstrate that the regret in linearized NSC is a specific instance within the OCO-M framework, achieved through a particular selection of separable functions (sub-section IV-A, visualized in Fig. 1 (up)). Second, we establish that the regret of OCO-M with these separable functions is in turn a particular case within the Delayed OCO (OCO-D) framework with a specific structure of delay (sub-section IV-B, visualized in Fig. 1 (down)). While the first part is fairly standard in NSC, we do not reduce its resulting OCO-M instance to standard OCO, but to OCO-D instead. The connection to delayed online learning was indeed identified in [19], but incorporating predictions (of unknown quality) was not considered.
@@
Given $d\geq\frac{1}{\delta}\log(\frac{z}{\delta\epsilon}T)$, $z\doteq w(\kappa_{B}\kappa_{M}p+1)$. Then, under Assumptions $1\&2$, for any $\epsilon>0$:
$\sum_{t=1}^{T}\bigg{|}c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})-\sum_{i=0}^{d}f_{t}^{(i)}(M_{t-i})\bigg{|}\leq\alpha\epsilon.$
@@
For now, assume that $\!\|\hat{\boldsymbol{x}_{t}}\!-\!\boldsymbol{x}_{t}\|\!\leq\!\frac{\epsilon}{T}$. Then, by Assump. 2:
$\sum_{t=1}^{T}|c_{t}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})-c_{t}(\boldsymbol{\hat{x}}_{t},\boldsymbol{u}_{t})|\leq\|\boldsymbol{\alpha}_{t}\|\ \|\hat{\boldsymbol{x}_{t}}\!-\!\boldsymbol{x}_{t}\|=\alpha\epsilon.$ (18)
However, $c_{t}(\boldsymbol{\hat{x}}_{t},\boldsymbol{u}_{t})$ can be written in terms of $f_{t}^{(i)}(\cdot)$ using again Assump. $2$ but with the definitions in (8) and (9) :
@@
It remains to show that $\|\boldsymbol{x}_{t}-\hat{\boldsymbol{x}_{t}}\|\leq\frac{\epsilon}{T}$. From (5) and (17):
$\|\hat{\boldsymbol{x}_{t}}-\boldsymbol{x}_{t}\|\leq\|\sum_{i=d}^{t-1}A^{i}\big{(}BM_{t-i-1}\overline{\boldsymbol{w}}_{t-i-2}+\boldsymbol{w}_{t-i-1}\!\big{)}\|$
$\leq\sum_{i=d}^{t-1}(\|A^{i}BM\|_{op}\|\boldsymbol{\overline{w}}_{t-i-2}\|+\|A^{i}\|_{op}\|\boldsymbol{w}_{t-i-1}\|)$
@@
where $(a)$ follow from the sub-multipilicitive property of $\|\cdot\|_{\text{op}}$, and $\|B\|_{\text{op}}\!\leq\!\|B\|$, $(b)$ by Assump. $1$, and the bounds on matrix norms, and $(c)$ from $1-x\leq e^{-x}$ and the definition of $z$. Substituting $d$, makes the last term $\frac{\epsilon}{T}$. ∎
Thus, the closeness of $\boldsymbol{x}_{t}$ and $\boldsymbol{\hat{x}}_{t}$ is due to the (strong) stability. Lastly, we note the possibility to set $d$ adaptively without knowledge of $T$ using $d=\frac{1}{\delta}\log(\frac{zt}{\delta\epsilon})$. [^3] This would result in a sum of the form $\sum_{t}\nicefrac{{1}}{{t}}\leq\log(T)$.
@@
Let $f_{t}(M_{0},\dots,M_{d})$ be a separable function: $f_{t}(M_{0},\dots,M_{d})=\sum_{i=0}^{d}f_{t,i}(M_{d-i})$, $f_{t,i}(M):\mathbb{R}^{(d_{u}\times d_{x}p)}\mapsto\mathbb{R}$ Let $\mathcal{A}$ be an online learning algorithm whose decisions $M_{t+1}$ depend on the history set $\mathbb{H}_{t}\doteq\cup_{i=0}^{d}\{f_{\tau,i}(\cdot)\}_{\tau=1}^{t}$. Define $J_{t}(M)\doteq\sum_{i=0}^{d}f_{t+i,i}(M)$, and the history set w.r.t. $J_{t}(\cdot)$ as $\mathbb{H}^{J}_{t}$ Then,
@@
(20) means that the accumulated cost, with memory, is equivalent to that of the memoryless functions $J_{t}(\cdot)$. However, from (19) $\mathcal{A}$ has delayed feedback w.r.t. $J_{t}(\cdot)$; when deciding $M_{t+1}$, feedback up to only $t-d$ is available.
@@
the first part is immediate from the definition of $J_{\tau}(\cdot)$; any $J_{\tau}(\cdot)$ with $\tau>t-d$ would require a function that is not in the original history set $\mathbb{H}_{t}$ (not known at $t$).
@@
Where the first equality holds by separability, the second by shifting the sum index and using the convention $f_{t<d}(\cdot)\doteq 0$ w.l.o.g, [^4] The adversarial rounds can always be prefixed with zero cost rounds. Alternatively, redefine regret to start from $t=d$ as in see [11, Sec 2.2]. and the third by using $f_{t>T}(\cdot)\doteq 0$.∎
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{t=1}^{T}\big{(}\sum_{i=0}^{d}f_{t}^{(i)}(M_{t-i})-\sum_{i=0}^{d}f_{t}^{(i)}(M^{\star})\big{)}+2\alpha\epsilon$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{t=1}^{T}f_{t}(M_{t-d},\dots,M_{t})-f_{t}(M^{\star},\dots,M^{\star})+2\alpha\epsilon$ (22)
[…]
where $(a)$ follows by Lemma 1, which gives both an upper bound on the learner cost and lower bound on the benchmark’s cost, $(b)$ by writing the sum of $f_{t}^{(i)}(\cdot)$ as a single function with memory, and $(c)$ by Lemma 2, with $f_{t,i}(M)=f_{t}^{(i)}(M)$, and hence, $J_{t}(M)=F_{t}(M)$. To bound the delayed feedback regret $\mathcal{R}_{T}^{F}$, we use [16, Thm. 11], which we restate below using the notation of this paper: [^5] Mapping their notation to ours, we get $\boldsymbol{g}_{t-D:t}=G_{t-d:t}$, $\boldsymbol{h}_{t}=H_{t}$, $\boldsymbol{g}_{t-D:t}-h_{t}=\Delta_{t}$, $\boldsymbol{a}_{t,F}=2\kappa_{M}\Delta_{t}$, $\boldsymbol{b}_{t,F}=\nicefrac{{1}}{{2}}\Delta^{2}_{t}$, and $\alpha=\kappa_{M}^{2}$.
@@
Theorem 2 ([16, Thm. 11])
Let $\lambda_{t}$ be non-decreasing on $t$ defined as in (16). Let $r_{t}(\cdot)$ be $\lambda_{t}$ strongly convex regularizer. Then, $\{M_{t}\}_{t}$ computed via (14) ensures that:
$\mathcal{R}_{T}^{F}\leq 8\kappa_{M}\max_{t\in[T]}\Delta_{t-d:t-1}+2\sqrt{5}\kappa_{M}\sqrt{\sum_{t=1}^{T}\Delta_{t}^{2}}.$ (24)
To write $\Delta_{t}$ explicitly, note that $G_{t-d:t}$ can be written as:
@@
[…]
$\sum_{i=0}^{d-1}\sum_{j=d-i}^{d}\|{G}^{(j)}_{t-d+i+j}-{\tilde{G}}^{(j)}_{t-d+i+j}\|+\sum_{j=0}^{d}\Delta^{(j)}_{t+j}=$ (26)
$\sum_{i=0}^{d}\sum_{j=d-i}^{d}\Delta_{t-d+i+j}^{(j)}=\!\sum_{i=0}^{d}\sum_{j=0}^{i}\Delta_{t+j}^{(j+d-i)}=\!\sum_{i=0}^{d}\sum_{j=i}^{d}\Delta_{t+i}^{(d-j+i)}$ (27)
Substituting back in (24) gives the result. ∎